# On compatibility of Koszul- and higher preprojective gradings

### v2

@@ 4. The $n$-representation finite case
Hence, $T$ becomes $\mathbb{Z}^{2}$-graded, and we refer to the cohomological grading and the internal grading respectively. Taking the degree 00 part of the internal grading, we obtain
@@
Since $R_{\bullet}$ is $(p,n+1)$-Koszul, by [BBK02, Proposition 3.2] the algebra $T$ is generated in cohomological degrees 00 and $1$, while the full Ext-algebra $\bigoplus_{i\geq 0}\operatorname{Ext}^{i}_{R}(R_{0},R_{0})$ is generated by its cohomological degree 00 and $1$ pieces together with $\operatorname{Ext}_{R}^{n+2}(R_{0},R_{0})$. However, since $A$ has global dimension $n$, it follows that $\operatorname{Ext}_{A}^{n+2}(R_{0},R_{0})=0$, and hence $\bigoplus_{i\geq 0}\operatorname{Ext}^{i}_{A}(R_{0},R_{0})$ is generated in cohomological degrees 00 and $1$. By 2.7, it follows that $A$ is a Koszul algebra. ∎
@@
Let $R_{\bullet}$ be $(p,n+1)$-Koszul such that there exists a higher preprojective grading ${}_{\bullet}R$, where $A={}_{0}R$ is $n$-representation finite. Then there exists a higher preprojective cut on $R_{\bullet}$ whose degree 00 component is isomorphic to $A$.
@@
Using 4.2, we obtain that $A$ is Koszul. Therefore, we can compute the $(n+1)$-total grading $\Pi_{n+1}(A)_{\bullet}$ on the higher preprojective algebra $\Pi_{n+1}(A)$. The higher preprojective grading then is a cut with respect to the $(n+1)$-total grading by 2.22. Note that $R\simeq\Pi_{n+1}(A)$ and that both rings are equipped with a $(p,n+1)$-Koszul grading. Since $R$ is finite dimensional, it follows that $R_{\bullet}\simeq\bigoplus_{i\geq 0}\operatorname{rad}(R)^{i}/\operatorname{rad}(R)^{i+1}$, and therefore $R_{\bullet}$ and $\Pi_{n+1}(A)_{\bullet}$ are graded isomorphic. Hence, transporting the higher preprojective grading on $\Pi_{n+1}(A)$ along this isomorphism, we obtain a higher preprojective cut on $R_{\bullet}$, and its degree 00 part is isomorphic to $A$. ∎
@@ 5.1. Graded radicals and Graded Simples
The purpose of this section is to discuss the graded simples of ${}_{\bullet}R$ and $R_{\bullet}$. We first show that ${}_{\bullet}R$ and $R_{\bullet}$ have isomorphic semisimple base rings in degree 00. However, we do not obtain that the graded simples coincide as ungraded modules, so we will need additional assumptions.
@@
Since both gradings are nonnegative, it follows that a graded simple module is concentrated in a single degree, where it is a simple module over the degree 00 part. That means a sum of representatives of the graded simple modules as ungraded modules is given by
@@ 5.2. The graded radicals
Note that $R$ is generated in degrees 00 and $1$ for both gradings. Furthermore, $\operatorname{grad}(R_{\bullet})$ is generated as an ideal by $R_{1}$, so if it is generated by a set of nilpotent elements, we obtain from 5.5 that this set lies in $\operatorname{rad}(A)\oplus{}_{+}R=\operatorname{grad}({}_{\bullet}R)$, and hence we have $\operatorname{grad}(R_{\bullet})\subseteq\operatorname{grad}({}_{\bullet}R)$. Similarly, $\operatorname{grad}({}_{\bullet}R)$ is generated as an ideal by $\operatorname{rad}(A)\oplus{}_{1}R$. Since $A$ is finite dimensional, $\operatorname{rad}(A)$ is generated by a set of nilpotent elements, and by assumption this extends to a generating set of nilpotent elements for $\operatorname{grad}({}_{\bullet}R)$, so 5.5 shows $\operatorname{grad}(R_{\bullet})\supseteq\operatorname{grad}({}_{\bullet}R)$. ∎
@@ 5.3. Nilpotency of generators
In this section, we assume that ${}_{\bullet}R$ is a higher preprojective grading so that ${}_{0}R$ is $n$-representation infinite. We show that the assumption on ${}_{1}R$ being generated by nilpotent elements is fulfilled automatically. We recall that we may interpret ${}_{1}R$ as
@@
so we analyse the structure of this space as an $A$-bimodule. In the following, we denote by $\Omega(M)$ a syzygy of the module $M$. Additionally, recall that in case $A$ is $n$-representation infinite one has the notion of n-preprojective modules — i.e. modules of the form $\tau^{-j}_{n}A$ — and note that such modules $M$ satisfy $\operatorname{Ext}_{A}^{i}(DA,M)=0$ for $i<n$.
@@
Let $A$ be $n$-hereditary for $n\geq 2$, and assume that $M$ is non-injective and satisfies $\operatorname{Ext}^{n-1}_{A}(DA,M)=0$. Then we have
$\Omega(\tau_{n}^{-}(M))\simeq\tau_{n-1}^{-}(M).$
@@
To compute $\tau_{n}^{-}(M)$, we choose a minimal injective coresolution
$I^{\bullet}\colon\quad 0\to M\to I^{0}\to I^{1}\to\cdots\to I^{n-1}\to I^{n}\to 0.$
Next, we apply the derived inverse Nakayama functor $\nu^{-1}$ and keep track of the part of the resulting complex that will be relevant.
$\cdots\to\nu^{-1}(I^{n-2})\to\nu^{-1}(I^{n-1})\to\nu^{-1}(I^{n})\to 0$
Taking the cokernel at the end gives us $\tau_{n}^{-}(M)$ by definition. We also note that $\nu^{-1}(I^{n})$ is projective and surjects onto $\tau_{n}^{-}(M)$. Furthermore, since we assumed in the beginning that the coresolution is minimal, the morphism $\nu^{-1}(I^{n})\to\tau_{n}^{-}(M)$ is a projective cover, hence the kernel is $\Omega(\tau_{n}^{-}(M))$.
${{\nu^{-1}(I^{n-2})}}$${{\nu^{-1}(I^{n-1})}}$${{\nu^{-1}(I^{n})}}$00{0}${{\Omega(\tau_{n}^{-}(M))}}$${{\tau_{n}^{-}(M)}}$$\scriptstyle{g}$$\scriptstyle{f}$
Next, we recall that $\tau_{n-1}^{-}(M)$ can be computed using the same resolution. To do this, take the $n-2$-th cosyzygy and apply $\tau^{-}$, which means we obtain $\tau_{n-1}^{-}(M)$ as the cokernel of $g$. Furthermore, since $f\circ g=0$, this induces an epimorphism $\tau_{n-1}^{-}(M)\to\Omega(\tau_{n}^{-}(M))$, and its kernel is precisely the homology of $\nu^{-1}(I^{\bullet})$ at position $n-1$.
${{\nu^{-1}(I^{n-2})}}$${{\nu^{-1}(I^{n-1})}}$${{\nu^{-1}(I^{n})}}$00{0}${{H_{n-1}(\nu^{-1}(I^{\bullet}))}}$${{\tau_{n-1}^{-}(M)}}$${{\Omega(\tau_{n}^{-}(M))}}$00{0}$\scriptstyle{g}$$\scriptstyle{f}$
Now, it suffices to note that $H_{n-1}(\nu^{-1}(I^{\bullet}))=\operatorname{Ext}_{A}^{n-1}(D(M),A)=0$ since $M$ is $n$-rigid. This gives us the desired isomorphism. ∎
@@
If $A$ is $n$-hereditary and $\operatorname{Hom}_{A}(D(A),A)=0$, we have that $\operatorname{Hom}_{A}(\Omega(\tau_{n}^{-}(P)),P)=0$ for any indecomposable projective $P$.
@@
From above, we have that $\Omega(\tau_{n}^{-}(P))\simeq\tau_{n-1}^{-}(P)$ holds, and so, using $\operatorname{Hom}_{A}(D(A),A)=0$, we get that
$\operatorname{Hom}_{A}(\Omega(\tau_{n}^{-}(P)),P)$ $\simeq\operatorname{Hom}_{A}(\tau_{n-1}^{-}(P),P)$
$\simeq\operatorname{Hom}_{A}(\tau^{-}(\Omega^{-(d-2)}(P)),P)$
$\simeq\operatorname{Ext}_{A}^{1}(P,\Omega^{-(d-2)}(P))$
$\simeq 0.\qed$
Proposition 5.9.
Let $A$ be $n$-hereditary, let $P$ be an indecomposable projective, and assume that $\operatorname{Hom}_{A}(D(A),A)=0$. Then any morphism from $P$ to $\tau_{n}^{-}(P)$ factors through some projective $A$-module $Q$ such that no summand of $Q$ is isomorphic to $P$.
@@
Since $P$ is projective, any morphism $P\to\tau_{n}^{-}(P)$ must factor through the projective cover $Q\to\tau_{n}^{-}(P)$. We show that $Q$ has no direct summand isomorphic to $P$. We know from 5.8 that $\operatorname{Hom}_{A}(\Omega(\tau_{n}^{-}(P)),P)=0$. Let’s now consider the short exact sequence
$0\to\Omega(\tau_{n}^{-}(P))\to Q\to\tau_{n}^{-}(P)\to 0$
and apply $\operatorname{Hom}_{A}(\--,P)$. This gives the beginning of a long exact sequence
[…]
Note that the term $\operatorname{Hom}_{A}(\tau_{n}^{-}(P),P)$ is zero since the higher preprojective component is directed, or equivalently because the higher preprojective algebra is nonnegatively graded. Furthermore, we see that the term $\operatorname{Ext}^{1}_{A}(\tau_{n}^{-}(P),P)$ is zero since we assumed that $n\geq 2$, and because the higher preprojective component is $n$-rigid. Thus, we find
$\operatorname{Hom}_{A}(Q,P)\simeq\operatorname{Hom}_{A}(\Omega(\tau_{n}^{-}(P)),P)=0,$
and therefore $Q$ has no summand isomorphic to $P$. ∎
Note that $\operatorname{Hom}_{A}(D(A),A)=0$ holds whenever $A$ is $n$-representation infinite. Let us now use this to prove the nilpotency we are interested in.
Proposition 5.10.
Let $A$ be $n$-representation infinite. Then there exists a set $E\subseteq\operatorname{Ext}^{d}_{A}(D(A),A)$ which generates $\Pi_{n+1}(A)$ as an $A$-algebra and which consists of nilpotent elements in $\operatorname{Ext}^{d}_{A}(D(A),A)\subseteq\Pi_{n+1}(A)$.
Proof.
We begin by recalling that
$\operatorname{Ext}^{d}_{A}(D(A),A)\simeq\operatorname{Hom}_{A}(A,\tau_{n}^{-}(A)).$
This space can be decomposed according to the decomposition of $A=\bigoplus_{i}P_{i}$ into indecomposable projectives, so we write
Clearly, any element $f\in\operatorname{Hom}_{A}(P_{i},\tau_{n}^{-}(P_{j}))$ for non-isomorphic $P_{i}\not\simeq P_{j}$ squares to 00 in the preprojective algebra, since that square is just the composition
Thus, it remains to deal with morphisms $f\in\operatorname{Hom}_{A}(P_{i},\tau_{n}^{-}(P_{i}))$. However, from 5.9, we know that $f$ factors through some projective $Q$ which has no summand isomorphic to $P_{i}$, and hence $f$ lies in the radical of the $A$-module $\bigoplus_{i,j}\operatorname{Hom}_{A}(P_{i},\tau_{n}^{-}(P_{j}))$, so we can ignore it for generating purposes by Nakayama’s lemma. ∎
@@ 5.4. Koszulity of $n$-representation infinite algebras
of $M$ by graded projective ${}_{\bullet}R$-modules $P^{i}$. This resolution is called minimal if $\operatorname{Im}(d_{i})\subseteq\operatorname{grad}(P^{i+1})$. Taking the degree 00 part of this resolutions then yields a minimal projective resolution of the ${}_{0}R$-module ${}_{0}M$ by projective ${}_{0}R$-modules.
@@
Since $S$ is concentrated in degree 00, it follows from the proof of [Mad11, Proposition 3.1.2] that
@@
for each $i\geq 0$. Since $R_{\bullet}$ is Koszul, we have that $\operatorname{Ext}^{0}_{R}(S,S)\simeq D(R_{0})$ and $\operatorname{Ext}^{1}_{R}(S,S)\simeq D(R_{1})$, which both are finite dimensional by assumption. Furthermore, since $R_{\bullet}$ is Koszul, we have that $\operatorname{Ext}^{\ast}_{R}(S,S)$ is generated in cohomological degrees 00 and $1$, so it follows that $\operatorname{Ext}_{R}^{i}(S,S)$ is finite dimensional for all $i\geq 0$. Therefore, each direct product $\prod_{j\in\mathbb{Z}}\operatorname{Ext}^{i}_{{}_{\bullet}R}(S,S\langle j\rangle)$ is a finite direct product, and hence we obtain
@@
is generated in cohomological degree 00 and $1$. It follows from 5.12 that
@@
is a subalgebra of $R^{!}=\operatorname{Ext}^{*}_{R}(S,S)$, and we claim that it is also generated in cohomological degree 00 and $1$. Assume to the contrary that it has a generator $x$ in cohomological degree $>1$. Since $x$ is also an element in $R^{!}$, it can be expressed as a sum of products of generators of $R^{!}$ that lie in degree 00 and $1$. However, as ${}_{\bullet}R$ is positively graded, we have that $R^{!}$ has no generators lying in some
@@
for $j<0$, which would be necessary if ${}_{0}R^{!}$ were to have a generator in cohomological degree $>1$, and hence there are no such generators $x$.
Now, we consider a minimal graded projective resolution $P^{\bullet}$ of $S$ over ${}_{\bullet}R$ as in Remark 5.11. Then its degree 00 part is a minimal projective resolution of $S$ over $A$. Observe that $\operatorname{grad}({}_{\bullet}R)$ annihilates $S$, and that thus we can write
@@
from which we deduce that $\operatorname{Ext}^{*}_{A}(S,S)$ is generated in degree 00 and $1$, and hence $A$ is Koszul by 2.7. ∎

### v3

@@ 4. The $n$-representation finite case
Hence, $T$ becomes $\mathbb{Z}^{2}$-graded, and we refer to the cohomological grading and the internal grading respectively. Taking the degree 0 part of the internal grading, we obtain
@@
Since $R_{\bullet}$ is $(p,n+1)$-Koszul, by [BBK02, Proposition 3.2] the algebra $T$ is generated in cohomological degrees 0 and $1$, while the full Ext-algebra $\bigoplus_{i\geq 0}\operatorname{Ext}^{i}_{R}(R_{0},R_{0})$ is generated by its cohomological degree 0 and $1$ pieces together with $\operatorname{Ext}_{R}^{n+2}(R_{0},R_{0})$. However, since $A$ has global dimension $n$, it follows that $\operatorname{Ext}_{A}^{n+2}(R_{0},R_{0})=0$, and hence $\bigoplus_{i\geq 0}\operatorname{Ext}^{i}_{A}(R_{0},R_{0})$ is generated in cohomological degrees 0 and $1$. By 2.7, it follows that $A$ is a Koszul algebra. ∎
@@
Let $R_{\bullet}$ be $(p,n+1)$-Koszul such that there exists a higher preprojective grading $_{\bullet}R$, where $A={}_{0}R$ is $n$-representation finite. Then there exists a higher preprojective cut on $R_{\bullet}$ whose degree 0 component is isomorphic to $A$.
@@
Using 4.2, we obtain that $A$ is Koszul. Therefore, we can compute the $(n+1)$-total grading $\Pi_{n+1}(A)_{\bullet}$ on the higher preprojective algebra $\Pi_{n+1}(A)$. The higher preprojective grading then is a cut with respect to the $(n+1)$-total grading by 2.22. Note that $R\simeq\Pi_{n+1}(A)$ and that both rings are equipped with a $(p,n+1)$-Koszul grading. Since $R$ is finite dimensional, it follows that $R_{\bullet}\simeq\bigoplus_{i\geq 0}\operatorname{rad}(R)^{i}/\operatorname{rad}(R)^{i+1}$, and therefore $R_{\bullet}$ and $\Pi_{n+1}(A)_{\bullet}$ are graded isomorphic. Hence, transporting the higher preprojective grading on $\Pi_{n+1}(A)$ along this isomorphism, we obtain a higher preprojective cut on $R_{\bullet}$, and its degree 0 part is isomorphic to $A$. ∎
@@ 5.1. Graded radicals and Graded Simples
The purpose of this section is to discuss the graded simples of $_{\bullet}R$ and $R_{\bullet}$. We first show that $_{\bullet}R$ and $R_{\bullet}$ have isomorphic semisimple base rings in degree 0. However, we do not obtain that the graded simples coincide as ungraded modules, so we will need additional assumptions.
@@
Since both gradings are nonnegative, it follows that a graded simple module is concentrated in a single degree, where it is a simple module over the degree 0 part. That means a sum of representatives of the graded simple modules as ungraded modules is given by
@@ 5.2. The graded radicals
Note that $R$ is generated in degrees 0 and $1$ for both gradings. Furthermore, $\operatorname{grad}(R_{\bullet})$ is generated as an ideal by $R_{1}$, so if it is generated by a set of nilpotent elements, we obtain from 5.5 that this set lies in $\operatorname{rad}(A)\oplus{}_{+}R=\operatorname{grad}({}_{\bullet}R)$, and hence we have $\operatorname{grad}(R_{\bullet})\subseteq\operatorname{grad}({}_{\bullet}R)$. Similarly, $\operatorname{grad}({}_{\bullet}R)$ is generated as an ideal by $\operatorname{rad}(A)\oplus{}_{1}R$. Since $A$ is finite dimensional, $\operatorname{rad}(A)$ is generated by a set of nilpotent elements, and by assumption this extends to a generating set of nilpotent elements for $\operatorname{grad}({}_{\bullet}R)$, so 5.5 shows $\operatorname{grad}(R_{\bullet})\supseteq\operatorname{grad}({}_{\bullet}R)$. ∎
@@ 5.3. Nilpotency of generators
In this section, we assume that $_{\bullet}R$ is a higher preprojective grading so that $_{0}R$ is $n$-representation infinite. We show that the assumption on $_{1}R$ being generated by nilpotent elements is fulfilled if we assume that the Gabriel quiver of $_{0}R=A$ is acyclic. The question whether the Gabriel quiver of an $n$-hereditary algebra is necessarily acyclic has been raised in [HIO14, Question 5.9] and remains, to the authors’ knowledge, open.
We recall that we may interpret $_{1}R$ as
@@
so we analyse the structure of this space as an $A$-bimodule. This space can be decomposed according to the decomposition of $A=\bigoplus_{i}P_{i}$ into indecomposable projectives, so we write
Clearly, any element $f\in\operatorname{Hom}_{A}(P_{i},\tau_{n}^{-}(P_{j}))$ for non-isomorphic $P_{i}\not\simeq P_{j}$ squares to 0 in the preprojective algebra, since that square is just the composition
Thus, it remains to deal with morphisms $f\in\operatorname{Hom}_{A}(P_{i},\tau_{n}^{-}(P_{i}))$.
Remark 5.7.
By [HIO14, Theorem 4.24], there exists an $n$-almost split sequence
$0\to P\xrightarrow[]{f_{n}}C_{n}\xrightarrow[]{f_{n-1}}\cdots\xrightarrow[]{f_{1}}C_{1}\xrightarrow[]{f_{0}}\tau_{n}^{-}(P)\to 0$
in the additive closure $\mathcal{P}=\operatorname{add}\{\tau_{n}^{-i}(A)\mid i\geq 0\}$. By [HIO14, Proposition 2.3 (b)] this additive closure is directed in the sense that
$\operatorname{Hom}_{\mathcal{P}}(\tau_{n}^{-j}(A),\tau_{n}^{-i}(A))=0$
for $j>i$. In the following, we write $C_{n+1}=P$ and $C_{0}=\tau_{n}^{-}(P)$ for convenience.
Thus, the terms $C_{i}$ in the $n$-almost split sequence consist of projective summands or iterated inverse translates of projectives. We first show that the appearance of a indecomposable projective summand $Q$ in $C_{i}$ means that there is a path in the Gabriel quiver of $A$ from the vertex supporting $P$ to the one supporting $Q$.
@@
Let $Q_{i}$ be an indecomposable projective summand of $C_{i}$ for some $1\leq i\leq n$. Then there exist indecomposable projective summands $Q_{j}$ of $C_{j}$ for $i<j\leq n+1$ such that every $f_{j-1}$ induces a non-zero morphism $Q_{j}\to Q_{j-1}$. Furthermore, $Q_{n+1}=P$.
@@
The proof is inductive in nature. We denote by $\pi_{i}\colon C_{i}\to Q_{i}$ the projection. We claim that $\pi_{i}\circ f_{i}\neq 0$. Indeed if $\pi_{i}\circ f_{i}=0$, consider the part of the $n$-almost split sequence
$C_{i+1}$$C_{i}$$C_{i-1}$$\scriptstyle{f_{i}}$$\scriptstyle{f_{i-1}}$
and apply $\operatorname{Hom}_{\mathcal{P}}(-,Q_{i})$. We obtain
$\operatorname{Hom}_{\mathcal{P}}(C_{i+1},Q_{i})$$\operatorname{Hom}_{\mathcal{P}}(C_{i},Q_{i})$$\operatorname{Hom}_{\mathcal{P}}(C_{i-1},Q_{i})$$\scriptstyle{-\circ f_{i}}$$\scriptstyle{-\circ f_{i}}$
Since the original diagram was part of an $n$-almost split sequence, the resulting diagram is exact. By assumption, we have that $\pi_{i}\in\operatorname{Hom}_{\mathcal{P}}(C_{i},Q_{i})$ is in $\operatorname{Ker}(-\circ f_{i})$. Hence by exactness, we have that $\pi_{i}\in\operatorname{Im}(-\circ f_{i-1})$. This means $\pi=g\circ f_{i-1}$ for some $g$, which contradicts the fact that $f_{i-1}$ is a radical morphism.
Thus, the composition $\pi_{i}\circ f_{i}$ is nonzero. Since $Q_{i}$ is projective, it follows that there exists a projective summand $Q_{i+1}$ of $C_{i+1}$ such that the restriction $(\pi_{i}\circ f_{i})|_{Q_{i+1}}$ is nonzero. Repeating the same argument for $Q_{i+1}$, we obtain a summand $Q_{i+2}$ of $C_{i+2}$, until we find a projective summand $Q_{n+1}$ of $C_{n+1}=P$, which has to be $Q_{n+1}=P$ since $P$ was assumed to be indecomposable. ∎
The following corollary is immediate.
@@
Using the same notation, if $P$ appears as a summand in some $C_{i}$ for $i<n+1$, then there is a cycle in the Gabriel quiver of $A$.
We now apply this to show that any morphism $P\to\tau_{n}^{-}(P)$ needs to factor in $\mathcal{P}$, unless $A$ has a cycle in its Gabriel quiver.
Proposition 5.10.
Let $f\colon P\to\tau_{n}^{-}(P)$ be an arbitrary morphism, and consider as before the $n$-almost split sequence
$0\to P\xrightarrow[]{f_{n}}C_{n}\xrightarrow[]{f_{n-1}}\cdots\xrightarrow[]{f_{1}}C_{1}\xrightarrow[]{f_{0}}\tau_{n}^{-}(P)\to 0.$
Then the following hold:
  1. (1)
    The morphism $f$ factors through $f_{0}$.
  2. (2)
    The module $C_{1}=Q_{1}\oplus\tau_{n}^{-}(Q_{1}^{\prime})$ for projective modules $Q_{1}$ and $Q_{1}^{\prime}$.
  3. (3)
    The module $P$ is not a direct summand of $Q_{1}^{\prime}$.
  4. (4)
    If $P$ is a direct summand of $Q_{1}$, then the Gabriel quiver of $A$ has a cycle.
@@
  1. (1)
    This is immediate because $f_{0}$ is right almost split.
  2. (2)
    By [HIO14, Lemma 4.25], the map $f_{0}$ is a sink map in $\mathcal{P}$. The statement then follows from $f_{0}$ being right minimal and $\operatorname{Hom}_{\mathcal{P}}(\tau_{n}^{-j}(A),\tau_{n}^{-}(P))=0$ for $j\geq 2$.
  3. (3)
    Suppose that $P$ was a summand of $Q_{1}^{\prime}$. Then $f_{0}$ restricted to $\tau_{n}^{-}(P)$ gives an irreducible morphism in $\operatorname{Hom}_{\mathcal{P}}(\tau_{n}^{-}(P),\tau_{n}^{-}(P))$. Applying $\tau_{n}$ then produces an irreducible morphism $g=\tau_{n}(f_{0}|_{\tau_{n}^{-}(P)})$ in $\operatorname{Hom}_{\mathcal{P}}(P,P)$. This means that $g$ is irreducible in $\operatorname{proj}(A)$, and so the Gabriel quiver of the finite dimensional algebra $A$ of finite global dimension has a loop, contradicting the no-loop theorem [ILP11].
  4. (4)
    This follows immediately by applying 5.9. ∎
We obtain the claimed nilpotency as a corollary for acyclic $n$-representation infinite algebras.
Corollary 5.11.
Let $A$ be an acyclic $n$-representation infinite algebra. Then the $A$-bimodule $\operatorname{Hom}_{A}(A,\tau_{n}^{-}(A))$ is generated by a set of elements that are nilpotent in the preprojective algebra. In particular, $\Pi_{n+1}(A)$ is generated as an algebra by $A$ and some nilpotent elements.
@@
We generate $\operatorname{Hom}_{A}(A,\tau_{n}^{-}(A))$ by generating $\operatorname{Hom}_{A}(P_{i},\tau_{n}^{-}(P_{j}))$ for every pair $(i,j)$. If $i\neq j$, all elements in $\operatorname{Hom}_{A}(P_{i},\tau_{n}^{-}(P_{j}))$ square to 0, and so we pick any basis of $\operatorname{Hom}_{A}(P_{i},\tau_{n}^{-}(P_{j}))$ and add it to our generating set. If $i=j$, then every element in $\operatorname{Hom}_{A}(P_{i},\tau_{n}^{-}(P_{i}))$ factors through $Q_{1}\oplus\tau_{n}^{-}(Q_{1})$ where $Q_{1}$ and $Q_{1}^{\prime}$ are projective modules without direct summands isomorphic to $P_{i}$. Hence, every element in $\operatorname{Hom}_{A}(P_{i},\tau_{n}^{-}(P_{i}))$ is generated by the previously chosen basis elements. ∎
@@ 5.4. Koszulity of $n$-representation infinite algebras
of $M$ by graded projective $_{\bullet}R-$modules $P^{i}$. This resolution is called minimal if $\operatorname{Im}(d_{i})\subseteq\operatorname{grad}(P^{i+1})$. Taking the degree 0 part of this resolutions then yields a minimal projective resolution of the $_{0}R-$module $_{0}M$ by projective $_{0}R-$modules.
@@
Since $S$ is concentrated in degree 0, it follows from the proof of [Mad11, Proposition 3.1.2] that
@@
for each $i\geq 0$. Since $R_{\bullet}$ is Koszul, we have that $\operatorname{Ext}^{0}_{R}(S,S)\simeq D(R_{0})$ and $\operatorname{Ext}^{1}_{R}(S,S)\simeq D(R_{1})$, which both are finite dimensional by assumption. Furthermore, since $R_{\bullet}$ is Koszul, we have that $\operatorname{Ext}^{\ast}_{R}(S,S)$ is generated in cohomological degrees 0 and $1$, so it follows that $\operatorname{Ext}_{R}^{i}(S,S)$ is finite dimensional for all $i\geq 0$. Therefore, each direct product $\prod_{j\in\mathbb{Z}}\operatorname{Ext}^{i}_{{}_{\bullet}R}(S,S\langle j\rangle)$ is a finite direct product, and hence we obtain
@@
is generated in cohomological degree 0 and $1$. It follows from 5.13 that
@@
is a subalgebra of $R^{!}=\operatorname{Ext}^{*}_{R}(S,S)$, and we claim that it is also generated in cohomological degree 0 and $1$. Assume to the contrary that it has a generator $x$ in cohomological degree >1>1. Since $x$ is also an element in $R^{!}$, it can be expressed as a sum of products of generators of $R^{!}$ that lie in degree 0 and $1$. However, as $_{\bullet}R$ is positively graded, we have that $R^{!}$ has no generators lying in some
@@
for $j<0$, which would be necessary if $_{0}R^{!}$ were to have a generator in cohomological degree >1>1, and hence there are no such generators $x$.
Now, we consider a minimal graded projective resolution $P^{\bullet}$ of $S$ over $_{\bullet}R$ as in Remark 5.12. Then its degree 0 part is a minimal projective resolution of $S$ over $A$. Observe that $\operatorname{grad}({}_{\bullet}R)$ annihilates $S$, and that thus we can write
@@
from which we deduce that $\operatorname{Ext}^{*}_{A}(S,S)$ is generated in degree 0 and $1$, and hence $A$ is Koszul by 2.7. ∎